# Quantformer: from attention to profit with a quantitative transformer trading strategy

### v2

@@ 5.3 Training under different scales
Understanding data volume and data frequency is critical to designing infrastructure for processing data and model stability [55]. To test the ability of stock selection of the model, three other factors ”QF_10%”, ”QF_5%”, ”QF_1%” are trained and tested (with ”QF_Month_1%, named as ”QF_20%”). The four groups are all trained under $d_{f}$ = 3, which include the first, middle, and tail 20%, 10%, 5%, and 1% of stocks in ranking, respectively. The backtest period and other settings are the same with 4.
@@
Table 4 shows the result of the backtest. QF_10strategy recorded an AR of 13.12% and an AER of 17.73%, with a slightly lower portfolio volatility of 0.159. This strategy had the highest WR at 61.8%. QF_5strategy delivered 12.59% AR and 16.02% AW, with the lowest portfolio volatility at 0.136. It had the lowest VaR at 2.015%. QF_1strategy outperformed the others, achieving the highest AR of 24.71% and the highest AER of 35.74%, though it had the highest volatility at 0.214. These four factors all perform better than CSI300 based on their return and risk. The result shows that the model takes the ability to select stocks under different training set and requirements.
@@ 6 Conclusion
This study proposed a new model, quantformer for quantitative stock prediction and trading. We addressed the need for handling numerical input data rather than text and adapting the model for forecasting tasks rather than sequence-to-sequence problems common in NLP. To enable direct processing of numerical time series data, we replaced the word embedding layer with a standard linear layer and removed the output masking operations. We also simplified the decoder to produce a probability distribution over future price movements rather than autoregressively generating token sequences.
Eventually, we tested the factor in the reality back test of the financial market and compared the factor with other 100 price-volume factors. Our experimental results demonstrate the promise of this approach. The model-based trading strategies were able to deliver substantial excess returns over the benchmark under a longer time period, with Sharpe Ratios indicating sound risk-adjusted performance. The positive alphas affirm the strategy’s ability to outperform expected returns after accounting for risk factors.
There is further scope to enhance the framework, for instance by incorporating additional signals like news and fundamentals as auxiliary inputs. The self-attention mechanism could likely be improved to encode greater temporal context, or use transformer-based models such as GPT-4 [56] and Claude3 [57] to establish the strategy as these models are fine-tuned and may perform better for quantitative finance tasks.
Overall, our work illustrates the viability of quantformer for financial data modeling, provides a flexible framework to model the market from multiple perspectives of financial time series and develops profitable trading strategies that consider different variables. With further research to solve the above limitations, we believe such ML-based quantitative methods hold rich potential. The framework provides a flexible foundation for better understanding markets and developing profitable trading strategies.
The implementation code of quantformer is available at https://github.com/zhangmordred/QuantFormer.
@@ 7 Acknowledgements and declaration
This work was funded by the Natural Science Foundation of China (12271047); Guangdong Provincial Key Laboratory of Interdisciplinary Research and Application for Data Science, BNU-HKBU United International College (2022B1212010006); UIC research grant (R0400001-22; UICR0400008-21; R72021114); Guangdong College Enhancement and Innovation Program (2021ZDZX1046).
@@ Appendix A Detailed backtest results between 100 factors and the quantformer factor
All the data of factors comes from JoinQuant [^4] https://www.joinquant.com/, which is one of the most common quantitative finance platforms. Detailed descriptions of the following factors are available at JoinQuant Factors.
@@
  • Annual Return (AR) and Annual Excess Return (AER) are two metrics to reflect the return of the portfolio directly. The excess returns are returns achieved above and beyond the return of a proxy and the CSI300 index is used as the basis when calculating the excess returns.
  • Win Rate (WR) is the metric to evaluate the fraction between the action that makes a profit and all the actions.
  • Sharpe Ratio (SR) is a measure of risk-adjusted return that describes the additional earnings an investor receives for each standard deviation unit increase, which is shown in equation (9).
  • The Alpha measures the ability of a portfolio to generate returns above the market benchmark. A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment, which is shown in equation (10).
  • Beta is the correlation between the portfolio and systemic risk, reflecting the sensitivity of the strategy to the change in the market. For the daily return of the strategy, $D_{p}$, and the daily return of the benchmark $D_{m}$, the equation of beta is:
    $Beta=\beta_{p}=\frac{Cov(D_{p},D_{m})}{Var(D_{m})}$
    where $Cov(D_{p},D_{m})$ is the correlation of between $D_{p}$ and $D_{m}$, $Var(D_{m})$ is the variance of daily return of benchmark.
  • Max Drawdown (MD) means the potential worst-case scenario or the most extreme possible loss from the previous peak. For the trough value of a portfolio $V_{trough}$ and the peak value of the portfolio $V_{peak}$, the MDD is calculated in the following way:
    $MDD=\frac{V_{peak}-V_{trough}}{V_{peak}}$
  • Portfolio volatility ($\sigma_{p}$) is the standard deviation of the portfolio returns. $\sigma_{p}$ avoids the problem of negative deviations canceling with positive deviations and also penalizes larger deviations from the mean. It provides a kind of weighted average deviation in which large deviations carry more weight [58].
@@ Appendix B Comparison between different quantformer strategies
(a) Factor performance
(b) Comparison between models
Figure 6: Comparison of factor performance and model performance
In the provided scatter plot 6(a), the $QF\_Month\_1$ factor exhibits notable performance in terms of annual return. Compared to other factors, $QF\_Month\_1$ offers better returns. Regarding risk, as measured by the $Variance\_Covariance\_VaR\_99$, the $QF\_Month\_1$ factor’s data points are positioned in a moderately high range, indicating a balanced risk profile. The size and color intensity of the $QF\_Month\_1$ points, reflecting higher Alpha values, suggest a robust excess return or relative performance. Combining these observations, the $QF\_Month\_1$ factor likely provides a compelling investment opportunity by delivering robust returns while maintaining a moderate risk level, an attractive proposition for strategies aiming to optimize the trade-off between risk and return.
In scatter plot 6(b), most of the factors based on quantformer perform better compared with the other 100 factors. The performance of quantformer factors which are trained under different scales shows ability in making profit and risk management.

### v3

@@ 5.2 Training under different scales
To test the stock selection capabilities of the model, three other factors “QF_10%”, “QF_5%”, “QF_1%” are trained and tested (with “QF_Month_1%”, named as “QF_20%”). The four groups are all trained under $\varrho$ = 3, which include the first, middle, and tail 20%, 10%, 5%, and 1% of stocks in ranking, respectively. The backtest period and other settings are the same as in Section 4.
@@
Table 4 shows the result of the backtest. QF_10% strategy records an AR of 13.12% and an AER of 17.73%, with a slightly lower portfolio volatility of 0.159. This strategy has the highest WR at 61.8%. QF_5% strategy delivered 12.59% AR and 16.02% AW, with the lowest portfolio volatility at 0.136. It has the lowest VaR at 2.015%. QF_1% strategy outperformed the others, achieving the highest AR of 24.71% and the highest AER of 35.74%, though it takes the highest volatility at 0.214. These four factors all perform better than CSI 300 based on their return and risk. These results suggest that the quantformer model’s performance seems robust to the choice of different training sets and requirements.
@@ 6 Conclusion
This study proposes a new neural network architecture, quantformer, inspired by the transformer architecture, for quantitative stock prediction and trading. We address the need for handling numerical input data rather than text, and adapted the model for forecasting tasks rather than sequence-to-sequence problems common in NLP. To enable direct processing of numerical time series data, we replace the word embedding layer with a standard linear layer and removed the output masking operations. We also simplify the decoder to produce a probability distribution over future price movements rather than autoregressively generating token sequences.
We also backtest the quantformer factor on financial market data and compare it with 100 other price-volume factors. Our experimental results demonstrate the promise of this approach. The quantformer-based trading strategies are able to deliver substantial excess returns over the benchmark under a longer time period, with Sharpe Ratios indicating sound risk-adjusted performance. The positive alphas affirm the strategy’s ability to outperform expected returns after accounting for risk factors.
Overall, our work illustrates the viability of the quantformer architecture to handle financial time series, design profitable and robust trading strategies, and calls for further analysis and investigations of this proposed ML-based quantitative methods. The implementation code of quantformer is available at https://github.com/zhangmordred/QuantFormer.
@@ 7 Acknowledgements and declaration
This work was funded by the Natural Science Foundation of China (12271047); Guangdong Provincial/Zhuhai Key Laboratory of Interdisciplinary Research and Application for Data Science, Beijing Normal-Hong Kong Baptist University (2022B1212010006); BNBU research grants (UICR0400008-21; UICR0700041-22; R72021114); and the Guangdong College Enhancement and Innovation Program (2021ZDZX1046).
@@ Appendix A Detailed backtest results between 100 factors and the quantformer factor
All the data of factors come from JoinQuant [^5] https://www.joinquant.com/, which is one of the most common quantitative finance platforms in China. Detailed descriptions of the following factors are available at https://www.joinquant.com/help/api/help#factor_values.
@@
  • Annual Return (AR) and Annual Excess Return (AER) are two metrics to reflect the return of the portfolio directly. The excess returns are returns achieved above and beyond the return of a proxy and the CSI 300 index is used as the basis when calculating excess returns.
  • Win Rate (WR) is the metric defined as the fraction between the number of trading periods that generate a profit over the total number of trading periods.
  • Sharpe Ratio (SR) is a measure of risk-adjusted return that describes the additional earnings an investor receives for each unit of the standard deviation of returns. It is shown in equation (17).
  • The Alpha measures the ability of a portfolio to generate returns above the market benchmark. A positive alpha indicates that the investment portfolio has achieved a higher return than its benchmark after risk adjustment. It is shown in equation (18).
  • Beta is the correlation between the portfolio and systemic risk, reflecting the sensitivity of the strategy to changes in the market. For the daily return of the strategy, $D_{p}$, and the daily return of the benchmark $D_{m}$, the equation of beta is:
    $\mathrm{Beta}=\beta_{p}=\frac{\mathbb{C}\mathrm{ov}(D_{p},D_{m})}{\mathbb{V}\hskip-0.84pt\mathrm{ar}(D_{m})}$
    where $\mathbb{C}\mathrm{ov}(D_{p},D_{m})$ is the correlation between between $D_{p}$ and $D_{m}$, and $\mathbb{V}\hskip-0.84pt\mathrm{ar}(D_{m})$ is the variance of the daily returns of the benchmark.
  • Max Drawdown (MD) means the potential worst-case scenario or the most extreme possible loss from the previous peak. For the trough value of a portfolio $A_{\mathrm{trough}}$ and the previous peak value of the portfolio $A_{\mathrm{peak}}$, the MD is calculated in the following way:
    $\mathrm{MD}=\frac{A_{\mathrm{peak}}-A_{\mathrm{trough}}}{A_{\mathrm{peak}}}$
  • Portfolio volatility ($\sigma_{p}$) is the standard deviation of the portfolio returns. It provides a kind of weighted average deviation in which large deviations carry more weight (Bacon, 2023).
@@ Appendix B Comparison between different quantformer strategies
Figure 5: Comparison between factor performance and quantformer performance
In the provided scatterplot 5, the $\mathrm{QF\_Month\_1}$ factor exhibits notable performance in terms of annual return. Compared to other factors, $\mathrm{QF\_Month\_1}$ offers better returns. Regarding risk, as measured by the
$\mathrm{Variance\_Covariance\_VaR\_99}$, the $\mathrm{QF\_Month\_1}$ factor’s data points are positioned in a moderately high range, indicating a balanced risk profile. The size and color intensity of the $\mathrm{QF\_Month\_1}$ points, reflecting higher Alpha values, suggest a robust excess return or relative performance. Combining these observations, the $\mathrm{QF\_Month\_1}$ factor likely provides a compelling investment opportunity by delivering robust returns while maintaining a moderate risk level, an attractive proposition for strategies aiming to optimize the trade-off between risk and return. Overall, most of the factors based on quantformer perform better compared with the other 100 factors.
Appendix C Symbol table
Symbol Description
$A_{\mathrm{peak}},A_{\mathrm{trough}}$ peak value of the portfolio
$\alpha$ alpha ratio
$\mathbf{b}$ decision factor in the strategy
$\theta$ bias in the model
$\beta_{p}$ beta ratio
$D_{p}$ daily return
$h,H$ attention head
$\chi^{t}_{n}$, $\tilde{\chi}^{t}_{n}$ input feature matrix, normalized sequence
$\hat{Y}^{t}_{n}$ predicted output
$\mathbb{E}[x]$ mean
$\mathcal{G}^{t}$ set of the next timestamp profit
$\mathrm{MD}$ max drawdown
$t$ timestep
$m$ the number of timesteps before $t$
$\mathrm{MSELoss}$ MSE loss
$n$ the index of the stock
$p^{t}$ the close price of the stock
$P$ amount value of portfolio
$\Phi$ trading strategy
$\Psi$ the empirical quantile CDF
$R_{f}$, $R_{m}$, $R_{p}$ return of risk-free rate, market return, return of portfolio
$r^{t}_{n}$ profit rate
$s^{t}_{n}$ stock
$S^{t}$ stock set
$\mathrm{SR}$ Sharpe ratio
$\mathrm{std}[x]$ standard deviation
$\mathrm{TR}$ turnover rate of the portfolio
$v^{t}_{n}$ turnover rate of the stock over the time period $t$
$\mathrm{VaR}_{\alpha}$ value at risk at $\alpha$ level
$w^{t}_{n}$ weight of portfolio
$x^{t}_{n}$, $\tilde{x}^{t}_{n}$ row vector, normalized vector
$y^{t}_{n}$, $\tilde{y}^{t}_{n}$ target label, sorted predicted output
$\varrho$ number of bins in $y^{t}_{n}$
$\varphi$ percentage of stocks for each $\varrho$
$\xi$ middle percentage of $y^{t}_{n}$
$\mathcal{L}$ loss of holding stock
$\mathbf{F}$ multi-head attention output
$\mathbf{K}$, $\mathbf{Q}$, $\mathbf{V}$ key, query, value in quantformer
$\mathbf{W}$ trainable weight matrix
$\mathbf{Z}$ output layer